# NJL-Chiral Soliton and the Nucleon Equation of State at supra-saturation density: Impact of Chiral Symmetry Restoration

### v1

@@ III.2.1 Scalar field dependency
To this end, we artificially vary the nuclear scalar field $s$, and compute the corresponding scalar field $\cal S$ using the relation provided in eq. 21. Given a fixed set of NJL parameters ($\{G_{1},G_{2},m_{q},\Lambda\}$) and the value of $\cal S$, we calculate the in-medium quantities: the pion decay constant $F_{\pi}({\cal S})$, pion mass $M_{\pi}({\cal S})$, vector meson mass $M_{v}({\cal S})$, vector coupling $g_{v}({\cal S})$, and the parameter $a({\cal S})$. With these medium-modified quantities, we solve for the soliton configuration following the numerical procedure described in section II.2.
@@
The local baryon density ($\rho_{B}(r)$) profiles (carried by valence quarks) of the soliton for various values of $s$ are shown in fig. 8. As the scalar field strength increases, which corresponds to more restored chiral symmetry (or higher matter density as discussed in section III.2.2), we observe that the quarks become progressively delocalized. This results in the baryon density spreading over a larger spatial region, effectively increasing the hard-core size of the nucleon as displayed in fig. 9. Simultaneously, the central density decreases due to the reduced effective spatial quark confinement. This behavior is also reflected in the energy density profiles, as illustrated in fig. 11.
For sufficiently large scalar fields $s$, the notion of a localized hard core begins to break down altogether. The evolution of the baryon charge radius (interpreted as the hard-core radius) and the isoscalar charge radius with respect to $s$ is presented in figs. 9 and 9, respectively. The in-medium effect on the size of the hard core, which increases with density, indicates that the available volume for the nucleons in a nuclear medium will decrease, and the overlap of the hard cores begins at a density smaller than that corresponding to the overlap density calculated from the vacuum baryon charge radius.
@@
To ensure consistency with the observed nucleon mass $M_{N}$ in vacuum ($s=0$), we initially scale the energy density $\varepsilon_{\rm slt}(r)$ by a normalization constant $\chi$ defined in eq. 39. For non-zero scalar field values corresponding to in-medium conditions, we retain the same scaling factor $\chi$ determined in vacuum. This ensures that the comparison across different values of $s$ reflects in-medium modifications to the nucleon mass $M^{*}_{N}=\chi M_{\rm slt}(s)$. Hence, we re-normalized the EoSs at each $s$ by the scaling factor $\chi$ evaluated in vacuum. The resulting in-medium soliton mass $M_{\rm slt}(s)$ as a function of the scalar field is shown in fig. 10, which demonstrates the continuous reduction of the soliton (nucleon) mass as chiral symmetry is progressively restored.
@@
The local soliton EoSs (both $p_{\rm{slt},\chi}(\varepsilon_{\rm{slt},\chi})$ and $p_{\rm{slt}}(\varepsilon_{\rm{slt}})$), obtained for various values of the scalar field $s$ are shown in fig. 11. In contrast to the argument of Ref. [58], where the effect of chiral symmetry restoration on the nucleon EoS was taken into account by rescaling the energy density and pressure by an in-medium pion decay constant $f^{*}_{\pi}$, to this end, we have demonstrated the effect of chiral symmetry restoration on the interior EoS of the soliton in a self-consistent manner. As the strength of the scalar field $s$ (or density) increases and reaches to $0.2f_{\pi}-0.3f_{\pi}$ (or equivalently with the in-medium pion decay constant ($F_{\pi}^{*}$) approaches 0.85-0.93 times the vacuum value $f_{\pi}$ as shown in  appendix C), the re-normalized (by the factor $\chi$) EoS reaches a stiffness compared to that of the other preferred models of the high density NS matter (as displayed for SLy4 and QHC18). Whereas, ignoring the scaling of the EoS by $\chi$, one would reach the stiffness of the NS EoSs roughly for $s\sim 0.5f_{\pi}$ equivalently $F_{\pi}^{*}\to 0.7f_{\pi}-0.75f_{\pi}$).
@@
Following the argument of Ref. [58], the in-medium soliton energy density $\varepsilon^{\rm Med-F}_{\rm slt}$ and pressure $p^{\rm Med-F}_{\rm slt}$ can be obtained by rescaling their vacuum counterparts as
@@
  • The in-medium soliton EoS for $F_{\pi}^{*}=0.8f_{\pi}$, obtained using eqs. 44 and 45 (following Ref. [58]), is shown by dashed lines in figs. 12(a) and 12(b).
@@
  • For comparison, in figs. 12(a) and 12(b), we also display the soliton EoS obtained in our framework for a scalar field ${\cal S}(s)$ such that $F_{\pi}^{*}({\cal S})=0.8f_{\pi}$ (see section II), shown by dot–dashed lines.
@@
We notice from fig. 12, that if one considers the rescaling approach of Ref. [58], as discussed in eqs. 44 and 45, it tends to overestimate the stiffness of the EoS compared to the fully self-consistent approach adopted in this work. Qualitatively, this difference can be understood as a result of the competing effects from the scalar and vector sectors. The decrease of $F_{\pi}$ with increasing $s$ tends to stiffen the EoS, while the simultaneous reduction of $M_{v}$ acts in the opposite direction, softening the EoS in the core region. This highlights that a simple rescaling prescription cannot capture the interplay between scalar and vector contributions.
@@
Figure 12: (a) Soliton pressure $p_{\rm slt}(r)$ versus energy density $\varepsilon_{\rm slt}(r)$ for the NJL-F parametrization. The vacuum EoS (blue solid line) is compared with the in-medium modified soliton EoS (red dashed line) obtained using eqs. 44 and 45 with $F^{*}_{\pi}=0.8f_{\pi}$, and with the self-consistent soliton EoS (dot-dashed line) from section II for a scalar field $\mathcal{S}$ such that $F^{*}_{\pi}(\mathcal{S})=0.8f_{\pi}$. (b) Same as (a), but for $p_{\rm slt,\chi}(r)$ versus $\varepsilon_{\rm slt,\chi}(r)$, with $\chi$ fixed to its vacuum value.
III.2.2 Density dependence
@@
In this section, we extend our investigation on the in-medium modification of soliton properties and the EoS induced by the restoration of chiral symmetry relating to dense nuclear matter. In our framework, we assume that quarks describe the vacuum dynamics through spontaneous chiral symmetry breaking and mesonic degrees of freedom, while nuclear matter itself is treated as a collection of nucleons interacting via the meson fields generated by the underlying quark dynamics. The nucleons (baryon number $1$ solitons) are therefore the relevant degrees of freedom for matter. A more microscopic description, based on solving the quark wavefunctions in the soliton background, could provide direct access to the finite spatial extent and thus allow a more consistent treatment of finite-size effects. In the present work, however, we neglect such corrections and consider instead that, within nuclear matter, quarks do not propagate explicitly and the Dirac sea contribution is subtracted, as it naturally appears in the chiral scalar potential of the underlying NJL model in its bosonized form. For simplicity, we further restrict our study to symmetric nuclear matter (SNM) and focus on the evolution of soliton properties in this environment.
@@
The energy density of nuclear matter, $\varepsilon_{N}$, at nuclear matter density ($\rho_{N}$) in the mean-field approximation is given by [94]:
@@
The in-medium, scalar-field–dependent effective nucleon mass is provided by the NJL–soliton model (see sections II.1 and III.2.1) and is expressed as
$M^{*}_{N}(s)=\chi\,M_{s}(s).$ (48)
@@
In the mean-field limit, the in-medium scalar field $s$ is obtained by minimizing the energy density eq. 46 with respect to $s$. This yields the equation of motion for the scalar field ‘$s$’:
@@
Although our starting formulation is self consistent embedding the $s$-dependent nucleon mass from the NJL–soliton model directly into the mean-field energy density—we find that due to the rapid increase of $g^{*}_{s}(s)$ within this framework, the solution of the in-medium gap equation eq. 51 ceases to exist beyond $\rho_{N}\sim 4\rho_{\rm sat}$. To enable exploration of soliton properties at higher densities, we therefore follow a common practice in RMF approaches and fix the scalar coupling to its vacuum value,
@@
Physically, confinement effects are also expected to moderate the abrupt drop of the soliton mass with $s$, supporting this approximation. Thus, for a given density $\rho_{N}$, the corresponding nuclear scalar field $s({\cal S})$ is obtained from the solution of eq. 51, after which the soliton properties are computed following section II.1. In this way, we investigate the impact of in-medium chiral symmetry restoration on soliton properties and the nuclear EoS.
@@
While the energy density inside each soliton core has a spatial profile, we implicitly assume that, on average, the quarks experience an effective energy density equal to $\varepsilon_{QC}$. Thus, at a given baryon density $\rho_{N}$ with the corresponding nuclear scalar field $s(\rho_{N})$, we determine the radial distance $r=r_{QC}$ from the center of a soliton core (obtained for the given $s(\rho_{N})$ following section II.1) at which the local energy density $\varepsilon_{\rm slt}(r_{QC},s(\rho_{N}))$ matches the bulk value $\varepsilon_{QC}$. That is,
@@
To ensure consistency when comparing the resulting EoS eq. 57 with the NS EoS models as shown in fig. 15, we rescale the energy density and pressure of eq. 57 using a vacuum renormalization factor $\chi$ to define,
@@
[…]
@@
When the interaction energy is taken into account, such that the quark degrees of freedom are readjusted to ensure that the $\tilde{\varepsilon}_{QC,\chi}$ coincides with the nuclear matter energy density $\varepsilon_{N}$ at the nuclear density $\rho_{N}$, i.e. $\tilde{\varepsilon}_{QC,\chi}\equiv\varepsilon_{N}$ as given in  eq. 46, a systematic shift in the EoS is observed. Specifically, this matching condition leads to a lower onset value of $\rho_{N}$ at which the soliton-based EoS ceases to be valid, thereby indicating an earlier transition to quark matter. This outcome stands in contrast to the approximation $\varepsilon_{QC,\chi}=M_{N}^{*}\rho_{N}$, where the interaction contribution is neglected, and which consequently predicts the termination of the soliton branch at higher densities. And given the energy density $\tilde{\varepsilon}_{QC,\chi}=\varepsilon_{N}$ at a density $\rho_{N}$, we define the corresponding pressure as $p_{\chi}(\varepsilon_{N})=p_{\rm slt,\chi}(\varepsilon_{N};s(\rho_{N}))$ and for instance this is demonstrated in fig. 14 for NJL-F parameter set. As one can notice with the re-definition of $\tilde{\varepsilon}_{QC,\chi}=\varepsilon_{N}$, the onset of the termination reduced to $6\rho_{\rm sat}$ from $9\rho_{\rm sat}$ resulting with the consideration of $\varepsilon_{QC,\chi}=M_{N}^{*}\rho_{N}$.
@@ IV Conclusions
We have further investigated the properties of NJL-solitons embedded in nuclear matter through a self-consistent framework that accounts for the effects of chiral symmetry restoration at finite density. By describing nuclear matter within the relativistic mean-field (RMF) formalism, where the mesonic interactions are derived from a bosonization approach consistent with the adopted soliton model from the underlying NJL model, we have explored the density-dependent modifications to soliton characteristics, such as charge radii, as well as their implications for the EoS. This approach allows us to explore the density-dependent EoS, revealing a gradual modification of nucleon structure in dense matter. The observed overlap of soliton cores and the corresponding breakdown of the soliton-based EoS at high densities point to the onset of hard-deconfinement at density $\sim 8\rho_{\rm sat}-10\rho_{\rm sat}$ depending upon the parameters, providing a physically transparent connection between in-medium chiral dynamics and the possible emergence of quark matter. The results presented in section III.2.2 are the artifact of the slow evolution of the scalar field with the density due to the low scalar couplings. We have also presented an alternative way to extract an EoS from the different energy-density and pressure profiles inside the nucleon by considering the central point $r=0$ obtained for different values of the scalar field $s$ fixed by the nuclear density.
@@
The soliton properties presented in this work can be further improved by means of several systematic extensions. In particular, one may consider the contributions of the axial meson in the NJL-soliton description, the effects of spin-isospin quantization, the role of the scalar field (whose slow spatial variation across the soliton has been neglected here), and the inclusion of an explicit confining potential. Extending the analysis beyond SNM to $\beta$-equilibrated neutron-star matter or pure neutron matter, as well as moving beyond the mean-field approximation toward a complete Hartree-Fock treatment, would also provide a more comprehensive understanding of the role of chiral symmetry restoration in the results presented in section III.2.2. Furthermore, alternative treatments of nuclear matter with explicit quark degrees of freedom and finite-size corrections [84, 95], such as simulating solitons in a Wigner-Seitz cell or employing a Skyrmion lattice crystal approach [88, 96, 97], may yield a different evolution of the scalar field and consequently lead to further refinements of the soliton properties and the resulting EoS.
Figure 17: $\sigma_{y}/y$ obtained using  eq. 60.
V Acknowledgment

### v2

@@ 3.2.1 Scalar field dependency
To this end, we artificially vary the nuclear scalar field $s$, and compute the corresponding scalar field $\cal S$ using the relation provided in Eq. (21). Given a fixed set of NJL parameters ($\{G_{1},G_{2},m_{q},\Lambda\}$) and the value of $\cal S$, we calculate the in-medium quantities: the pion decay constant $F_{\pi}({\cal S})$, pion mass $M_{\pi}({\cal S})$, vector meson mass $M_{v}({\cal S})$, vector coupling $g_{v}({\cal S})$, and the parameter $a({\cal S})$. With these medium-modified quantities, we solve for the soliton configuration following the numerical procedure described in Section 2.2.
@@
The local baryon density ($\rho_{B}(r)$) profiles (carried by valence quarks) of the soliton for various values of $s$ are shown in Fig. 8. As the scalar field strength increases, which corresponds to more restored chiral symmetry (or higher matter density as discussed in Section 3.2.2), we observe that the quarks become progressively delocalized. This results in the baryon density spreading over a larger spatial region, effectively increasing the hard-core size of the nucleon as displayed in Fig. 9. Simultaneously, the central density decreases due to the reduced effective spatial quark confinement. This behavior is also reflected in the energy density profiles, as illustrated in Fig. 11.
For sufficiently large scalar fields $s$, the notion of a localized hard core begins to break down altogether. The evolution of the baryon charge radius (interpreted as the hard-core radius) and the isoscalar charge radius with respect to $s$ is presented in Figs. 9, 9, respectively. The in-medium effect on the size of the hard core, which increases with density, indicates that the available volume for the nucleons in a nuclear medium will decrease, and the overlap of the hard cores begins at a density smaller than that corresponding to the overlap density calculated from the vacuum baryon charge radius.
@@
To ensure consistency with the observed nucleon mass $M_{N}$ in vacuum ($s=0$), we initially scale the energy density $\varepsilon_{\rm slt}(r)$ by a normalization constant $\chi$ defined in Eq. (39). For non-zero scalar field values corresponding to in-medium conditions, we retain the same scaling factor $\chi$ determined in vacuum. This ensures that the comparison across different values of $s$ reflects in-medium modifications to the nucleon mass $M^{*}_{N}=\chi M_{\rm slt}(s)$. Hence, we re-normalized the EoSs at each $s$ by the scaling factor $\chi$ evaluated in vacuum. The resulting in-medium soliton mass $M_{\rm slt}(s)$ as a function of the scalar field is shown in Fig. 10, which demonstrates the continuous reduction of the soliton (nucleon) mass as chiral symmetry is progressively restored.
@@
The local soliton EoSs (both $p_{\rm{slt},\chi}(\varepsilon_{\rm{slt},\chi})$ and $p_{\rm{slt}}(\varepsilon_{\rm{slt}})$), obtained for various values of the scalar field $s$ are shown in Fig. 11. In contrast to the argument of Ref. Fukushima2020 , where the effect of chiral symmetry restoration on the nucleon EoS was taken into account by rescaling the energy density and pressure by an in-medium pion decay constant $f^{*}_{\pi}$, to this end, we have demonstrated the effect of chiral symmetry restoration on the interior EoS of the soliton in a self-consistent manner. As the strength of the scalar field $s$ (or density) increases and reaches to $0.2f_{\pi}-0.3f_{\pi}$ (or equivalently with the in-medium pion decay constant ($F_{\pi}^{*}$) approaches 0.85-0.93 times the vacuum value $f_{\pi}$ as shown in   C), the re-normalized (by the factor $\chi$) EoS reaches a stiffness compared to that of the other preferred models of the high density NS matter (as displayed for SLy4 and QHC18). Whereas, ignoring the scaling of the EoS by $\chi$, one would reach the stiffness of the NS EoSs roughly for $s\sim 0.5f_{\pi}$ equivalently $F_{\pi}^{*}\to 0.7f_{\pi}-0.75f_{\pi}$).
@@
Following the argument of Ref. Fukushima2020 , the in-medium soliton energy density $\varepsilon^{\rm Med-F}_{\rm slt}$ and pressure $p^{\rm Med-F}_{\rm slt}$ can be obtained by rescaling their vacuum counterparts as
@@
  • The in-medium soliton EoS for $F_{\pi}^{*}=0.8f_{\pi}$, obtained using Eqs. (44) and (45) (following Ref. Fukushima2020 ), is shown by dashed lines in Figs. 12(a) and  12(b), respectively.
@@
  • For comparison, in Figs. 12(a), and  12(b), we also display the soliton EoS obtained in our framework for a scalar field ${\cal S}(s)$ such that $F_{\pi}^{*}({\cal S})=0.8f_{\pi}$ (see Section 2), shown by dot–dashed lines.
@@
We notice from Fig. 12, that if one considers the rescaling approach of Ref. Fukushima2020 , as discussed in Eqs. (44) and (45), it tends to overestimate the stiffness of the EoS compared to the fully self-consistent approach adopted in this work. Qualitatively, this difference can be understood as a result of the competing effects from the scalar and vector sectors. The decrease of $F_{\pi}$ with increasing $s$ tends to stiffen the EoS, while the simultaneous reduction of $M_{v}$ acts in the opposite direction, softening the EoS in the core region. This highlights that a simple rescaling prescription cannot capture the interplay between scalar and vector contributions.
@@
Figure 12: (a) Soliton pressure $p_{\rm slt}(r)$ versus energy density $\varepsilon_{\rm slt}(r)$ for the NJL-F parametrization. The vacuum EoS (blue solid line) is compared with the in-medium modified soliton EoS (red dashed line) obtained using Eqs. (44)-(45) with $F^{*}_{\pi}=0.8f_{\pi}$, and with the self-consistent soliton EoS (dot-dashed line) from Section 2 for a scalar field $\mathcal{S}$ such that $F^{*}_{\pi}(\mathcal{S})=0.8f_{\pi}$. (b) Same as (a), but for $p_{\rm slt,\chi}(r)$ versus $\varepsilon_{\rm slt,\chi}(r)$, with $\chi$ fixed to its vacuum value.
3.2.2 Density dependence
@@
In this section, we extend our investigation on the in-medium modification of soliton properties and the EoS induced by the restoration of chiral symmetry relating to dense nuclear matter. In our framework, we assume that quarks describe the vacuum dynamics through spontaneous chiral symmetry breaking and mesonic degrees of freedom, while nuclear matter itself is treated as a collection of nucleons interacting via the meson fields generated by the underlying quark dynamics. The nucleons (baryon number $1$ solitons) are therefore the relevant degrees of freedom for matter. In the present work, however, we neglect such corrections and consider instead that, within nuclear matter, quarks do not propagate explicitly and the Dirac sea contribution is subtracted, as it naturally appears in the chiral scalar potential of the underlying NJL model in its bosonized form. Though the three-momentum cut-off on the underlying NJL model is not covariant, its use is standard and commonly adopted in NJL-based approaches Klevansky1992 ; Fukushima2013 . A more microscopic description, based on solving the quark wavefunctions in the soliton background, could provide direct access to the finite spatial extent and thus allow a more consistent treatment of finite-size effects. For simplicity, we further restrict our study to symmetric nuclear matter (SNM) and focus on the evolution of soliton properties in this environment.
@@
The energy density of nuclear matter, $\varepsilon_{N}$, at nuclear matter density ($\rho_{N}$) in the mean-field approximation is given by Chanfray2010b :
@@
The in-medium, scalar-field–dependent effective nucleon mass is provided by the NJL–soliton model (see Sections 2.1 and 3.2.1) and is expressed as
$M^{*}_{N}(s)=\chi\,M_{\rm slt}(s).$ (48)
@@
In the mean-field limit, the in-medium scalar field $s$ is obtained by minimizing the energy density Eq. (46) with respect to $s$. This yields the equation of motion for the scalar field ‘$s$’:
@@
Although our starting formulation is self consistent embedding the $s$-dependent soliton mass obtained from the NJL-soliton model directly into the mean-field energy density—we find that, when the density-dependent scalar coupling $g_{s}^{*}(s)$ (defined in Eq. (52)) is considered, the solution of the in-medium gap equation Eq. (51) ceases to exist beyond $\rho_{N}\sim 3\rho_{\rm sat}$. This breakdown originates from the rapid growth of $g_{s}^{*}(s)$ with increasing density, which causes the source term $g_{s}^{*}(s)\,n_{S}$ to increase too strongly to be balanced by the scalar potential term $\partial V/\partial s$. As a consequence, the scalar field equation no longer admits a physical solution, preventing a meaningful exploration of soliton properties at higher densities within this fully self consistent setup. To enable exploration of soliton properties at higher densities, we therefore follow a common practice in RMF approaches and fix the scalar coupling to its vacuum value, as given in Eq. (54).
@@
Figure 13: Upper panel: The density dependency of the nuclear scalar field ($s$) as a function of the density for different NJL parameters considered in this work. The solutions are obtained by solving Eq. (51) with fixing $g_{s}^{*}$ from  Eq. (54). Lower Panel: same as upper panel but for the scalar field $\cal S$, obtained using Eq. (21).
Physically, confinement effects are also expected to moderate the abrupt drop of the soliton mass with $s$, supporting this approximation. Thus, for a given density $\rho_{N}$, the corresponding nuclear scalar field $s({\cal S})$ is obtained from the solution of Eq. (51) and the density evolution of $\cal S$} is displayed in Fig. 13. Then at a given ${\cal S}(\rho_{N})$ the soliton properties are computed following Section 2.1. In this way, we investigate the impact of in-medium chiral symmetry restoration on soliton properties and the nuclear EoS.
@@
While the energy density inside each soliton core has a spatial profile, we implicitly assume that, on average, the quarks experience an effective energy density equal to $\varepsilon_{QC}$. Thus, at a given baryon density $\rho_{N}$ with the corresponding nuclear scalar field $s(\rho_{N})$, we determine the radial distance $r=r_{QC}$ from the center of a soliton core (obtained for the given $s(\rho_{N})$ following Section 2.1) at which the local energy density $\varepsilon_{\rm slt}(r_{QC},s(\rho_{N}))$ matches the bulk value $\varepsilon_{QC}$. That is,
@@
To ensure consistency when comparing the resulting EoS Eq. (57) with the NS EoS models as shown in Fig. 16, we rescale the energy density and pressure of Eq. (57) using a vacuum renormalization factor $\chi$ to define,
@@
With this procedure, a density-dependent EoS is constructed directly from microscopic soliton profiles. At low baryon densities $\rho_{N}$, the small bulk energy density $\varepsilon_{QC}$ corresponds to soliton configurations at large distances $r_{QC}$, where the local pressure becomes negative, as illustrated in Fig. 16. In Fig. 17, we display the evolution of the squared speed of sound ($c^{2}_{s}=\frac{dp}{d\varepsilon}$) for different EoSs of Fig. 16. The results of Fig. 16 and Fig. 17 demonstrate that a consistent treatment of chiral symmetry restoration produces a stiff soliton-based EoS (i.e, $p_{QC,\chi}(\varepsilon_{QC,\chi})$) compared to the vacuum soliton EoS (i.e, $p_{\rm slt,\chi}(\varepsilon_{\rm slt,\chi})$) and in quantitative agreement with established high-density NS models.
@@
When the interaction energy is taken into account, such that the quark degrees of freedom are readjusted to ensure that the $\tilde{\varepsilon}_{QC,\chi}$ coincides with the nuclear matter energy density $\varepsilon_{N}$ at the nuclear density $\rho_{N}$, i.e. $\tilde{\varepsilon}_{QC,\chi}\equiv\varepsilon_{N}$ as given in  Eq. (46), a systematic shift in the EoS is observed. Specifically, this matching condition leads to a lower onset value of $\rho_{N}$ at which the soliton-based EoS ceases to be valid, thereby indicating an earlier transition to quark matter. This outcome stands in contrast to the approximation $\varepsilon_{QC,\chi}=M_{N}^{*}\rho_{N}$, where the interaction contribution is neglected, and which consequently predicts the termination of the soliton branch at higher densities. And given the energy density $\tilde{\varepsilon}_{QC,\chi}=\varepsilon_{N}$ at a density $\rho_{N}$, we define the corresponding pressure as $p_{\chi}(\varepsilon_{N})=p_{\rm slt,\chi}(\varepsilon_{N};s(\rho_{N}))$ and for instance this is demonstrated in Fig. 15 for NJL-F parameter set. As one can notice with the re-definition of $\tilde{\varepsilon}_{QC,\chi}=\varepsilon_{N}$, the onset of the termination reduced to $6\rho_{\rm sat}$ from $9\rho_{\rm sat}$ resulting with the consideration of $\varepsilon_{QC,\chi}=M_{N}^{*}\rho_{N}$.
@@ 4 Conclusions
Within the mean-field treatment of symmetric nuclear matter in Section 3.2.2, the effect of the surrounding medium on the soliton is encoded through uniform background fields, allowing the $\omega$ meson to asymptotically approach its in-medium mean-field value via a modified boundary condition. Implementing this prescription for the $\omega$ field in dense matter, in the same spirit as discussed in  Li2024 , we find that there is no significant change in the core properties of the soliton such as the soliton based EoS, hard-core radius. We find that with this modification, neither the qualitative behavior nor the quantitative results of our analysis are altered (see  D). For completeness, we have also examined an alternative implementation inspired by the Wigner–Seitz (WS) construction Nagai2006 , in which the soliton is embedded in a finite spherical cell whose radius is fixed by the baryon density and the boundary conditions are imposed at the cell boundary. As discussed in  D, this scenario does not lead to stable soliton solution within the present framework and therefore requires a more detailed and systematic investigation, which we defer to future work.
@@
We have further investigated the properties of NJL-solitons embedded in nuclear matter through a self-consistent framework that accounts for the effects of chiral symmetry restoration at finite density. By describing nuclear matter within the relativistic mean-field (RMF) formalism, where the mesonic interactions are derived from a bosonization approach consistent with the adopted soliton model from the underlying NJL model, we have explored the density-dependent modifications to soliton characteristics, such as charge radii, as well as their implications for the EoS. This approach allows us to explore the density-dependent EoS, revealing a gradual modification of nucleon structure in dense matter. The observed overlap of soliton cores and the corresponding breakdown of the soliton-based EoS at high densities point to the onset of hard-deconfinement at density $\sim 8\rho_{\rm sat}-10\rho_{\rm sat}$ depending upon the parameters, providing a physically transparent connection between in-medium chiral dynamics and the possible emergence of quark matter. The results presented in Section 3.2.2 are the artifact of the slow evolution of the scalar field with the density due to the low scalar couplings. We have also presented an alternative way to extract an EoS from the different energy-density and pressure profiles inside the nucleon by considering the central point $r=0$ obtained for different values of the scalar field $s$ fixed by the nuclear density.
@@
The soliton properties presented in this work can be further improved by means of several systematic extensions. In particular, one may consider the contributions of the axial meson in the NJL-soliton description, the effects of spin-isospin quantization, the role of the scalar field (whose slow spatial variation across the soliton has been neglected here), and the inclusion of an explicit confining potential. Extending the analysis beyond SNM to $\beta$-equilibrated neutron-star matter or pure neutron matter, as well as moving beyond the mean-field approximation toward a complete Hartree-Fock treatment, would also provide a more comprehensive understanding of the role of chiral symmetry restoration in the results presented inSection 3.2.2. Furthermore, alternative treatments of nuclear matter with explicit quark degrees of freedom and finite-size corrections Bentz2025 ; JAMINON1989 , such as simulating solitons in a Wigner-Seitz cell or employing a Skyrmion lattice crystal approach Ma2013 ; Ma2014 ; Dong2013 , may yield a different evolution of the scalar field and consequently lead to further refinements of the soliton properties and the resulting EoS.
Figure 18: $\sigma_{y}/y$ obtained using  Eq. (60).
5 Acknowledgment